# references.bib

### v1

@@ IV Model Evaluation and Results
As the field of 3D reconstruction is growing rapidly, we select the following representative set of object- and scene-level reconstruction methods. For a short description for each selected model, see Appendix A.
@@
SF3D [boss_sf3d_2024], InstantMesh [li_instant3d_2023], One2345 [liu_one-2-3-45_2023], LGM [tang_lgm_2025], Michelangelo [zhao_michelangelo_2023], ZeroShape [huang_zeroshape_2024], Real3D [jiang_real3d_2024], and DSO [li_dso_2025].
@@
MIDI [huang_midi_2024], Gen3DSR [dogaru_generalizable_2024], and PhysGen3D [chen_physgen3d_2025].
Since we are interested in estimating the Lagrangian state from visual observation, we do not consider methods that generate 3D structures only from text input (e.g. [chen_fantasia3d_2023, poole_dreamfusion_2022, chen_atlas3d_2024]). Furthermore, given our focus on single-view reconstruction, we do not evaluate models that require multiple viewpoints (e.g. [chen_2l3_2024, wang_pf-lrm_2023, xu_grm_2025, yu_monosdf_2022, wu_object-compositional_2022, li_rico_2023, kong_vmap_2023]) or even dense observations (including [chabal_online_2025, ni_phyrecon_2024, turkulainen_dn-splatter_2025, jiang_vr-gs_2024, patel_real--sim--real_2024, jia_discoverse_2024, barcellona_dream_2024, abou-chakra_physically_2024, han_re3sim_2025, torne_reconciling_2024, wu_rl-gsbridge_2024, lou_robo-gs_2024, li_robogsim_2024, pfaff_scalable_2025, qureshi_splatsim_2024, zhu_vr-robo_2025, jiang_vr-gs_2024]). Most established simulation engines in robotics require mesh representations of objects. At the same time, converting 3D Gaussians to mesh surfaces is not a trivial operation [huang_2d_2024]. To fairly assess reconstruction performance, we opt to only consider methods that either produce meshes themselves or come with pre-defined mesh extraction routines. Further, optimisation-based routines such as score distillation are slow by design. Since we are concerned with computational requirements and responsiveness during reconstruction (Section II), we do not consider such methods in our experiments (including [liu_zero-1--3_2023, liu_syncdreamer_2023, qian_magic123_2023, sun_dreamcraft3d_2023, tang_dreamgaussian_2023, chen_atlas3d_2024, shi_mvdream_2023, voleti_sv3d_2025, kim_complete_2025, wu_reconfusion_2024]). Finally, while we generally outline potential issues with offloading compute from the robot platform to cloud services in Section II, we are practically limited to evaluating methods that have open-source implementations available.
@@ IV-A Evaluation Data
To properly evaluate reconstruction methods, we require richly annotated real-world data. However, most scene-level datasets for reconstruction or SLAM only contain 2.5D information [silberman_indoor_2012, dai_scannet_2017, chang_matterport3d_2017]. We select two datasets in our experiments: parts of the YCB-Video dataset [xiang_posecnn_2018] and the Project Aria Digital Twin dataset [pan_aria_2023]. These datasets are particularly representative for robotics applications as they capture different aspects of the challenges faced in real-world manipulation tasks. The YCB-Video dataset features standardised objects with varying geometric complexity that have become benchmarks in the robotic manipulation community, while the Aria dataset contains everyday household items arranged in naturalistic configurations that robots would encounter in domestic environments. Together, they span a spectrum from controlled, well-defined object arrangements to complex, cluttered real-world scenes. Both datasets provide ground-truth meshes for all objects, their poses in each camera frame, and additional information such as segmentation masks and camera intrinsics of their respective camera types. This rich annotation is crucial for quantitatively evaluating reconstruction accuracy against true object geometry. See Figure 16 for example inputs from both datasets. For both datasets, we manually select objects and frames to ensure coverage of different object types, camera view-points, and object poses.
IV-B Experimental Methodology
Each reconstruction model is evaluated on the same set of objects and camera frames. As all evaluated methods are single-view reconstruction approaches, we provide them with a single image of the object to be reconstructed. For each method, we follow the provided data preparation routing, including cropping and resizing of the object to a common size in the image frame. For all methods that rely on off-the-shelf segmentation models in their pipeline, we provide ground-truth object masks during the reconstruction process. This is done to ensure a fair and consistent comparison independent from off-the-shelf methods.
To evaluate reconstruction error, we need to align the reconstructed mesh to the ground truth. Past works [liu_one-2-3-45_2023, li_dso_2025] have used alignment procedures that we extend as follows. First, we estimate the correct scale of the object. To do this, we compute the standard deviations of the mesh vertices along the top three principle components of each mesh. We then evaluate the ratio of the component-aligned standard deviations of the ground truth mesh w.r.t. the reconstruction and estimate the rescaling parameter as the median of these principal component-aligned ratios.
@@ IV-B Experimental Methodology
In our evaluations, we mainly rely on the Chamfer distance (CD) as a measure of reconstruction error. This is in line with recent reconstruction works [liu_zero-1--3_2023, xu_instantmesh_2024, hong_lrm_2023, li_dso_2025]. We follow the implementation of [mescheder_occupancy_2019] for computing the Chamfer distances.
@@ IV-C Results
SF3D [boss_sf3d_2024]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
@@
Figure 1: Reconstruction error on the YCB-Video [xiang_posecnn_2018] dataset as measured by the Chamfer distance (CD). CD is averaged across 10,000 sampled surface points from both target and reconstruction. Indicators for min, max, and median are shown. For most methods, reconstructed surfaces are about 5mm distant from the closest target mesh surface.
SF3D [boss_sf3d_2024]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
00{0}2020{20}204040{40}406060{60}608080{80}80100100{100}100Success Rate (%)3293282613461918224531389723319
Figure 2: Success rates of transferring grasp poses computed on the reconstruction to the target mesh. Indicated are 95%-confidence Wilson intervals and the number of successfully sampled grasps. Evaluated on the YCB-Video [xiang_posecnn_2018] dataset. The object reconstructions of which the Chamfer distances are reported in Figure 2 are too inaccurate to successfully transfer grasp poses from reconstruction to the target meshes. As a result, grasps that are computed and evaluated in simulation do not normally transfer to the real world.
SF3D [boss_sf3d_2024]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
@@
Figure 3: Total number of objects with physically stable poses within 5 from the ground truth scene pose versus total number of objects without such stable poses. Evaluated on YCB-Video. For results on the Aria Digital Twin [pan_aria_2023] dataset, see Figure 14. Object mesh reconstructions from all methods exhibit significant physical instability, with many objects unable to maintain proper positioning in physics simulations. Consequently, these scene reconstructions would collapse or disintegrate when subjected to physical simulation environments.
@@
To evaluate how accurately popular reconstruction models produce 3D reconstructions when queried with imagery that reflects realistic robotic observations, we run all models on segmented objects that are entirely visible except for self-occlusion. We select 177 such objects from YCB-Video and 136 objects from the Aria dataset. As there are no guarantees that reconstructed meshes are valid volumes, we evaluate reconstruction accuracy by uniformly sampling 10,000 points from the surfaces of ground truth mesh and reconstruction, and computing the Chamfer distance ($\ell^{2}$-norm) between both point clouds.
@@
Figures 2 and 13 depict the resulting Chamfer distances for the single-object models on the two datasets. On YCB-Video [xiang_posecnn_2018], with the exception of SF3D [boss_sf3d_2024] and DSO [li_dso_2025], reconstructed surfaces tend to be about 5mm distant from the closest surface in the ground truth meshes. Reconstruction accuracy on the Meta Aria Digital Twin [pan_aria_2023] dataset is significantly worse, with about double the median reconstruction error. Figure 16 depicts a collection of example scenes alongside mesh reconstructions to give the reader an impression of the visual manifestation of such error ranges. Models appear to especially struggle with objects in non-canonical poses, such as view-points towards the top or bottom of the object. Such reconstruction error would result in very different behaviour than the ground truth object, if put into a physics simulator. Further, we evaluate the transfer success rates of grasp poses computed on reconstructions to the ground truth meshes on YCB-Video [xiang_posecnn_2018]. The results in Figure 2 indicate that the reconstructed meshes are insufficiently accurate for computing grasp poses on the reconstructed meshes and applying them in the real world scene -- only about a third of all grasp poses can successfully be applied to the target meshes. For grasp sampling, we use the antipodal grasp sampling procedure presented in [gilles_metagraspnet_2022] with a Franka Emika Panda parallel gripper model. Grasp poses computed on the reconstructions are considered to transfer successfully if they are not in collision with the ground truth mesh, and the rotations between gripper surface normals and mesh surface normals at the grasp points are not in excess of 22.5, as specified in the implementation of [gilles_metagraspnet_2022].
@@
To evaluate the reconstruction quality of models that process entire scenes, we select 28 scenes from the YCB-Video [xiang_posecnn_2018] dataset, sample 10,000 points each from the reconstructed and ground truth surfaces of the entire scene, align the reconstructed scene as a whole with the target scene using the ICP process described in Section IV-B, and compute the Chamfer distances between reconstruction and target, as we have done in the single-object case. Figure 6 shows that models that reconstruct entire scenes perform worse on average than single-object reconstruction models. With the median error being between 1cm and 1.5cm, the reconstructed scene and its objects would behave very differently when used in a physics simulation. Such error ranges are too large for robust robotic manipulation and miss our formulated reconstruction target of 2mm by a large margin. Figure 11 separately evaluates the pose estimation of scene reconstruction models by looking at the relative Chamfer distance between two objects in the reconstruction and computing the difference to the relative distance between the same objects in the ground truth scene. The results show that the relative object poses are not preserved in the reconstruction, which is a requirement for both accurate physics simulation and being able to transfer manipulation trajectories from simulation to the real world. Figure 17 depicts several reconstruction examples of the scene-level models. It is clear that both reconstruction accuracy of individual objects as well as pose estimation is suboptimal.
YCB [xiang_posecnn_2018]
SF3D [boss_sf3d_2024]
Aria [pan_aria_2023]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
@@
Figure 4: Frequency of reconstructed scenes where reconstructed meshes are in collision with one another. Objects being more frequently in collision than not results in unwanted behaviour in physics simulation, such as instantaneous acceleration and collapse of the scene composition. As a result, these reconstructions are not suitable for use in robotics simulation. As the reconstruction models in this figure only operate on individual objects, this may well be outside the scope of these models. As we show in Figure 12, scene-level reconstruction models tend to produce fewer collisions.
0.00.51.01.52.06.0Chamfer Distance (cm)MIDI [huang_midi_2024]Gen3DSR [dogaru_generalizable_2024]PhysGen3D [chen_physgen3d_2025]
Figure 5: Reconstruction error of scene reconstruction models on the YCB-Video [xiang_posecnn_2018] dataset as measured by the Chamfer distance (CD). CD is averaged across 10,000 sampled surface points from the entire scene reconstruction and target. Indicators for min, max, and median are shown. Reconstructed surfaces tend to be over 1cm distant from the closest target surface.
50100200500700Reconstruction Time per Sample (s)  MIDI [huang_midi_2024]Gen3DSR [dogaru_generalizable_2024]PhysGen3D [chen_physgen3d_2025]051015202530Peak VRAM Allocation (GB)
Figure 6: Reconstruction time in seconds and memory utilisation for reconstructing entire scenes with multiple objects using different scene reconstruction models. While VRAM utilisation is roughly in line with what is reported for single-object models in Figure 11, the reconstruction time is an order of magnitude greater when operating on entire scenes. This effectively prohibits real-time execution when utilising such models.
050100120Number of MeshesMIDI [huang_midi_2024]Gen3DSR [dogaru_generalizable_2024]PhysGen3D [chen_physgen3d_2025]StableUnstable
@@
050100120Number of MeshesMIDI [huang_midi_2024]Gen3DSR [dogaru_generalizable_2024]PhysGen3D [chen_physgen3d_2025]StableUnstable
Figure 8: Object stability as in Figure 8 but where each reconstructed object mesh is separately aligned with the target mesh, instead of relying on the relative object poses from the scene reconstruction model. Individual alignment results in a greater number of stable poses, indicating that the scene reconstruction models produce suboptimal object poses.
@@
To evaluate how often inaccurate reconstructions result in mesh collisions, we select object pairs that are spatially close, reconstruct them individually, place them at their respective scene poses obtained through the ICP procedure described in Section IV-B, and use FCL [pan_fcl_2012] to compute mesh collisions between them. The evaluation dataset consists of 96 samples for YCB-Video [xiang_posecnn_2018] and 64 samples for the Meta Aria Digital Twin [pan_aria_2023] dataset.
Figure 4 shows that all single-object reconstruction models produce meshes that are in collision in a majority of the evaluation scenes. If used in a physics simulation, such collisions at initialisation would result in unwanted behaviour, such as instantaneous acceleration and collapse of the scene composition. Figure 12 shows that even though scene reconstruction models operate on entire scenes and thus in principle should be in a good position to avoid mesh collisions in their reconstructions through conditioning one object’s reconstructed mesh on other objects’ scene occupancy, as well as having control over the positioning of individual objects, their ability to avoid mesh collisions is only on par with the better single-object models. Only PhysGen3D shows vastly reduced mesh collisions. However, this may be an artifact of the model significantly overestimating relative object distances (Figure 11) which would automatically result in reduced object collisions.
@@
To evaluate whether reconstructed objects would be stable when put back into the scene, we first compute their stable orientations, that are are all orientations where the centre of mass is within the 2D convex hull of vertices that have minimal elevation. We then filter for those stable orientations where the z-axis is at most 5 tilted from the z-axis of the object in the scene, i.e. where the stable pose is similarly upright as the object in the scene. For each of the resulting candidate poses, we generate 8 equally spaced perturbations that rotate the object 5 away from its stable pose. If in a subsequent physics simulation using PyBullet [coumans_pybullet_2016] at least one of these perturbed poses leads to the object reverting back to the unperturbed pose, the unperturbed pose is considered stable. As a result, we obtain all poses that are physically robust to slight perturbation while also being close to the ground truth pose of the object in the scene. In our evaluation, we only consider objects that are resting on planes, not ones that are leaning against other objects. We use 146 samples from YCB-Video [xiang_posecnn_2018] and 133 samples from the Meta Aria Digital Twin [pan_aria_2023] dataset.
Figures 3 shows that the vast majority of YCB-Video reconstructions does not have stable poses near the observed scene pose. Michelangelo [zhao_michelangelo_2023] stands out among all other evaluated models with two thirds of the reconstructed meshes having valid stable poses. On the Meta Aria Digital Twin [pan_aria_2023] dataset, however, Michelangelo does not show the same advantage. Instead, all evaluated models produce even fewer stable mesh reconstructions than on YCB-Video. As a result, the reconstructed meshes would be largely unusable for physics simulation without significant, non-trivial postprocessing, as without it the object meshes would topple over as soon as physics steps are advanced.
Figure 8 shows that the scene reconstruction models do not fare any better. Instead, PhysGen3D [chen_physgen3d_2025] shows the least number of stable meshes among all evaluated models. To evaluate whether the reconstructed meshes themselves or the pose estimation of the scene reconstruction models are to blame, we align each reconstructed object mesh with the ground truth mesh separately and compute the number of stable poses again. The results are shown in Figure 8. The number of stable poses increases significantly, corroborating that the scene reconstruction models struggle to estimate the correct object poses. Still, the number of stable poses remains too low for practical use in robotics applications.
@@
SF3D [boss_sf3d_2024]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
@@
Scene reconstruction models fare much better in comparison. Figure 15 shows that across all models, the increase in median Chamfer distance is at most about 10%, with both MIDI [huang_midi_2024] and PhysGen3D [chen_physgen3d_2025] being significantly below that threshold. MIDI considers objects jointly during reconstruction, and PhysGen3D and Gen3DSR both utilise image inpainting models to complete partial occlusions before mesh reconstruction. Both approaches appear well-suited for maintaining reconstruction accuracy at occluded object regions. For this evaluation, we individually aligned each reconstructed object to the target object mesh to avoid compounding error from erroneous pose estimations by the model, as shown in Figure 11.
-0.50-0.250.000.250.500.751.001020Chamfer Distance Error (m)[huang_midi_2024][dogaru_generalizable_2024][chen_physgen3d_2025]
Figure 10: Relative object position error of scene-level reconstruction models PhysGen3D [chen_physgen3d_2025], Gen3DSR [dogaru_generalizable_2024], and MIDI [huang_midi_2024]. Indicators for min, max, median, and 25th and 75th quantile are shown. We evaluate the difference between the Chamfer distances (CD) between all object pairs in the ground truth scene, and the CDs between the corresponding object pairs in the reconstructed scene with relative object poses estimated by the reconstruction models. Scene reconstruction models tend to overestimate relative object distances, resulting in inaccurate scene reconstructions and potentially impaired scene stability.
SF3D [boss_sf3d_2024]
InstantMesh [xu_instantmesh_2024]
One2345 [liu_one-2-3-45_2023]
LGM [tang_lgm_2025]
Michelangelo [zhao_michelangelo_2023]
ZeroShape [huang_zeroshape_2024]
Real3D [jiang_real3d_2024]
DSO [li_dso_2025]
0.5151050100Reconstruction Time per Sample (s)  00{0}55{5}51010{10}101515{15}152020{20}202525{25}25Peak VRAM Allocation (GB)

### v2

@@ IV Model Evaluation and Results
As the field of 3D reconstruction is growing rapidly, we select the following representative set of object- and scene-level reconstruction methods.
@@
SF3D [80] extends TripoSR [68] by adding material and lighting estimation. Uses a transformer to predict a Triplane [201], which is meshed via differentiable Marching Tetrahedron [202]. InstantMesh [123] uses a multi-view diffusion model to generate views, which are encoded into Triplane [201] features. These features are then converted into meshes. One2345 [26] produces dense multi-view images via diffusion, then aggregates them into a discretized volume. This is converted into a signed distance function and finally meshed. LGM [96] generates multi-view images with diffusion and processes them via U-Net [131] into 3D Gaussians. These are rendered to train a NeRF [203], which is used for mesh extraction. Michelangelo [28] combines image, shape, and text in a latent autoencoder, followed by diffusion-based shape sampling. A conditioned occupancy network then extracts meshes. ZeroShape [142] estimates depth and camera intrinsics to project the image into 3D space. A conditional occupancy network processes this for mesh generation. Real3D [71] builds on TripoSR [68] and incorporates real-world data using an unsupervised loss. This co-training reduces domain shift and improves performance on real data. DSO [190] fine-tunes TRELLIS [204] with a physics-informed reward system for stable 3D shapes. TRELLIS encodes images into structured latents, decoded into SDFs for mesh extraction.
@@
MIDI [154] segments all objects in the scene using an image segmentation model. Then denoises latent embeddings jointly across objects using cross-instance self-attention, improving spatial reasoning. This enhances the model’s understanding of object relationships in the scene. Gen3DSR [153] segments and inpaints occluded objects before applying single-object 3D reconstruction. Aligns meshes to the scene using RANSAC [205] between object meshes and estimated depth. Background reconstruction is omitted for fair comparison in experiments. PhysGen3D [196] segments, inpaints, and reconstructs objects with InstantMesh [27]. Aligns them using keypoint descriptors and a differentiable renderer for refined pose. Adds language-model-inferred physical properties but does not handle collisions or stability during reconstruction.
Since we are interested in estimating the complete scene geometry from visual observation, we do not consider methods that generate 3D structures only from text input (e.g. [95, 50, 74]). Furthermore, given our focus on single-view reconstruction, we do not evaluate models that require multiple viewpoints (e.g. [128, 144, 100, 114, 156, 78, 77]) or even dense observations (including [206, 72, 151, 207, 24, 22, 15, 181, 16, 23, 19, 18, 17, 25, 20, 21, 207]). Most established simulation engines in robotics require mesh representations of objects. At the same time, converting 3D Gaussians to mesh surfaces is not a trivial operation [104]. To fairly assess reconstruction performance, we opt to only consider methods that either produce meshes themselves or come with pre-defined mesh extraction routines. Further, optimisation-based routines such as score distillation are slow by design. Since we are concerned with computational requirements and responsiveness during reconstruction (Section II), we do not consider such methods in our experiments (including [70, 69, 79, 188, 189, 74, 127, 198, 208, 209]). Finally, while we generally outline potential issues with offloading compute during deployment from the robot platform to cloud services in Section II, our empirical evaluation is practically constrained to methods with publicly available open-source implementations.
@@ IV-A Evaluation Data
We require richly annotated real-world data with complete 3D information. For our experiments,we select two datasets that are representative for robotics applications: parts of the YCB-Video dataset [38] and the Project Aria Digital Twin dataset [39]. The YCB-Video dataset features standardised objects with varying geometric complexity that have become benchmarks in the robotic manipulation community, while the Aria dataset contains everyday household items arranged in naturalistic configurations that robots would encounter in domestic environments. Together, they span a spectrum from controlled, well-defined object arrangements to complex, cluttered real-world scenes. Both datasets provide ground-truth meshes for all objects, their poses in each camera frame, and additional information such as segmentation masks and camera intrinsics of their respective camera types. This rich annotation is crucial for quantitatively evaluating reconstruction accuracy against true object geometry. See Fig. 17 for example inputs from both datasets. For both datasets, we manually select objects and frames to ensure coverage of different object types, camera view-points, and object poses.
IV-B Experimental Setup and Common Preprocessing
Each model received identical single-view inputs. For each method, we follow the provided data preparation routing, including cropping and resizing of the object to a common size in the image frame. For all methods that rely on off-the-shelf segmentation models in their pipeline, we provide ground-truth object masks during the reconstruction process. This is done to ensure a fair and consistent comparison independent from off-the-shelf methods.
To evaluate reconstruction error, we need to align the reconstructed mesh to the ground truth. Past works [26, 190] have used alignment procedures that we extend as follows. First, we estimate the correct scale of the object. To do this, we compute the standard deviations of the mesh vertices along the top three principle components of each mesh. We then evaluate the ratio of the component-aligned standard deviations of the ground truth mesh w.r.t. the reconstruction and estimate the rescaling parameter as the median of these principal component-aligned ratios.
@@ IV-C Results
SF3D [80]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
@@
Figure 2: Chamfer distances on YCB-Video [38]. CD is averaged across 10,000 sampled surface points from both target and reconstruction. Indicators for min, max, and median are shown. Most methods achieve  5mm surface.
SF3D [80]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
0{0}20{20}40{40}60{60}80{80}100{100}Success Rate (%)311343171248602646209377
Figure 3: Grasp transfer success rates from reconstruction to target mesh on YCB-Video [38]. Indicated are 95%-confidence Wilson intervals and the number of successfully sampled grasps. Low success rates indicate insufficient accuracy for manipulation tasks.
SF3D [80]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
@@
Figure 4: Total number of objects with physically stable poses within 5 from the ground truth scene pose versus total number of objects without such stable poses. Evaluated on YCB-Video. For results on the Aria Digital Twin [39] dataset, see Fig. 15. Object mesh reconstructions from all methods exhibit significant physical instability, with many objects unable to maintain proper positioning in physics simulations. Consequently, these scene reconstructions would collapse or disintegrate when subjected to physical simulation environments.
@@
To evaluate reconstruction accuracy on realistic robotic imagery, we run all models on segmented objects that are entirely visible except for self-occlusion. We select 177 such objects from YCB-Video and 136 objects from the Aria dataset. As there are no guarantees that reconstructed meshes are valid volumes, we evaluate reconstruction accuracy by uniformly sampling 10,000 points from the surfaces of ground truth mesh and reconstruction, and computing the Chamfer distance ($\ell^{2}$-norm) between both point clouds. Utilisation of the Chamfer distance (CD) as a measure of reconstruction error is in line with recent reconstruction works [70, 27, 135, 190]. We follow the implementation of [107] for computing the Chamfer distances.
@@
Figs. 3 and 14 depict the resulting Chamfer distances for the single-object models on the two datasets. On YCB-Video [38], with the exception of SF3D [80] and DSO [190], reconstructed surfaces tend to be about 5mm distant from the closest surface in the ground truth meshes. Reconstruction accuracy on the Aria Digital Twin [39] dataset is significantly worse, with about double the median reconstruction error. Fig. 17 depicts a collection of example scenes alongside mesh reconstructions to give the reader an impression of the visual manifestation of such error ranges. Models appear to especially struggle with objects in non-canonical poses, such as view-points towards the top or bottom of the object. Such reconstruction error would result in very different behaviour than the ground truth object, if put into a physics simulator. Further, we evaluate the transfer success rates of grasp poses computed on reconstructions to the ground truth meshes on YCB-Video [38]. The results in Fig. 3 indicate that the reconstructed meshes are insufficiently accurate for computing grasp poses on the reconstructed meshes and applying them in the real world scene - only about half of all grasp poses can successfully be applied to the target meshes. For grasp sampling, we use the antipodal grasp sampling procedure presented in [210] with a Franka Emika Panda parallel gripper model. Grasp poses computed on the reconstructions are considered to transfer successfully if they are not in collision with the ground truth mesh, the rotations between gripper surface normals and mesh surface normals at the grasp points are not in excess of 22.5, and the object is still in contact with the gripper after grasping and shaking it in simulation, as specified in the implementation of [210]. [210] demonstrate their grasp sampling procedure on common household objects, including YCB objects that also feature in our evaluation.
@@
To evaluate the reconstruction quality of models that process entire scenes, we select 28 scenes from the YCB-Video [38] dataset, sample 10,000 points each from the reconstructed and ground truth surfaces of the entire scene, align the reconstructed scene as a whole with the target scene using the ICP process described in Section IV-B, and compute the Chamfer distances between reconstruction and target, as we have done in the single-object case. Fig. 7 shows that models that reconstruct entire scenes perform worse on average than single-object reconstruction models. With the median error being between 1cm and 1.5cm, the reconstructed scene and its objects would behave very differently when used in a physics simulation. Such error ranges are too large for robust robotic manipulation and miss our formulated reconstruction target of 2mm by a large margin. Fig. 12 separately evaluates the pose estimation of scene reconstruction models by looking at the relative Chamfer distance between two objects in the reconstruction and computing the difference to the relative distance between the same objects in the ground truth scene. The results show that the relative object poses are not preserved in the reconstruction, which is a requirement for both accurate physics simulation and being able to transfer manipulation trajectories from simulation to the real world. Fig. 18 depicts several reconstruction examples of the scene-level models. It is clear that both reconstruction accuracy of individual objects as well as pose estimation is suboptimal.
YCB [38]
SF3D [80]
Aria [39]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
@@
Figure 5: Frequency of reconstructed scenes where reconstructed meshes are in collision with one another. Objects being more frequently in collision than not results in unwanted behaviour in physics simulation, such as instantaneous acceleration and collapse of the scene composition. As a result, these reconstructions are not suitable for use in robotics simulation. As the reconstruction models in this figure only operate on individual objects, this may well be outside the scope of these models. As we show in Fig. 13, scene-level reconstruction models tend to produce fewer collisions.
0.00.51.01.52.06.0Chamfer Distance (cm)MIDI [154]Gen3DSR [153]PhysGen3D [196]
Figure 6: Reconstruction error of scene reconstruction models on the YCB-Video [38] dataset as measured by the Chamfer distance (CD). CD is averaged across 10,000 sampled surface points from the entire scene reconstruction and target. Indicators for min, max, and median are shown. Reconstructed surfaces tend to be over 1cm distant from the closest target surface.
50100200500700Reconstruction Time per Sample (s)  MIDI [154]Gen3DSR [153]PhysGen3D [196]051015202530Peak VRAM Allocation (GB)
Figure 7: Reconstruction time in seconds and memory utilisation for reconstructing entire scenes with multiple objects using different scene reconstruction models. While VRAM utilisation is roughly in line with what is reported for single-object models in Fig. 12, the reconstruction time is an order of magnitude greater when operating on entire scenes. This effectively prohibits real-time execution when utilising such models.
050100120Number of MeshesMIDI [154]Gen3DSR [153]PhysGen3D [196]StableUnstable
@@
050100120Number of MeshesMIDI [154]Gen3DSR [153]PhysGen3D [196]StableUnstable
Figure 9: Object stability as in Fig. 9 but where each reconstructed object mesh is separately aligned with the target mesh, instead of relying on the relative object poses from the scene reconstruction model. Individual alignment results in a greater number of stable poses, indicating that the scene reconstruction models produce suboptimal object poses.
@@
To evaluate how often inaccurate reconstructions result in mesh collisions, we select object pairs that are spatially close, reconstruct them individually, place them at their respective scene poses obtained through the ICP procedure described in Section IV-B, and use FCL [211] to compute mesh collisions between them. The evaluation dataset consists of 96 samples for YCB-Video [38] and 64 samples for the Aria Digital Twin [39] dataset.
Fig. 5 shows that all single-object reconstruction models produce meshes that are in collision in a majority of the evaluation scenes. If used in a physics simulation, such collisions at initialisation would result in unwanted behaviour, such as instantaneous acceleration and collapse of the scene composition. Fig. 13 shows that even though scene reconstruction models operate on entire scenes and thus in principle should be in a good position to avoid mesh collisions in their reconstructions through conditioning one object’s reconstructed mesh on other objects’ scene occupancy, as well as having control over the positioning of individual objects, their ability to avoid mesh collisions is only on par with the better single-object models. Only PhysGen3D shows vastly reduced mesh collisions. However, this may be an artifact of the model significantly overestimating relative object distances (Fig. 12) which would automatically result in reduced object collisions.
@@
To evaluate whether reconstructed objects would be stable when put back into the scene, we first compute their stable orientations, that are are all orientations where the centre of mass is within the 2D convex hull of vertices that have minimal elevation. We then filter for those stable orientations where the z-axis is at most 5 tilted from the z-axis of the object in the scene, i.e. where the stable pose is similarly upright as the object in the scene. For each of the resulting candidate poses, we generate 8 equally spaced perturbations that rotate the object 5 away from its stable pose. If in a subsequent physics simulation using PyBullet [41] at least one of these perturbed poses leads to the object reverting back to the unperturbed pose, the unperturbed pose is considered stable. As a result, we obtain all poses that are physically robust to slight perturbation while also being close to the ground truth pose of the object in the scene. In our evaluation, we only consider objects that are resting on planes, not ones that are leaning against other objects. We use 146 samples from YCB-Video [38] and 133 samples from the Aria Digital Twin [39] dataset.
Fig. 4 shows that the vast majority of YCB-Video reconstructions does not have stable poses near the observed scene pose. Michelangelo [28] stands out among all other evaluated models with two thirds of the reconstructed meshes having valid stable poses. On the Aria Digital Twin [39] dataset, however, Michelangelo does not show the same advantage. Instead, all evaluated models produce even fewer stable mesh reconstructions than on YCB-Video. As a result, the reconstructed meshes would be largely unusable for physics simulation without significant, non-trivial postprocessing, as without it the object meshes would topple over as soon as physics steps are advanced.
Fig. 9 shows that the scene reconstruction models do not fare any better. Instead, PhysGen3D [196] shows the least number of stable meshes among all evaluated models. To evaluate whether the reconstructed meshes themselves or the pose estimation of the scene reconstruction models are to blame, we align each reconstructed object mesh with the ground truth mesh separately and compute the number of stable poses again. The results are shown in Fig. 9. The number of stable poses increases significantly, corroborating that the scene reconstruction models struggle to estimate the correct object poses. Still, the number of stable poses remains too low for practical use in robotics applications.
@@
SF3D [80]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
@@
Scene reconstruction models fare much better in comparison. Fig. 16 shows that across all models, the increase in median Chamfer distance is at most about 10%, with both MIDI [154] and PhysGen3D [196] being significantly below that threshold. MIDI considers objects jointly during reconstruction, and PhysGen3D and Gen3DSR both utilise image inpainting models to complete partial occlusions before mesh reconstruction. Both approaches appear well-suited for maintaining reconstruction accuracy at occluded object regions. For this evaluation, we individually aligned each reconstructed object to the target object mesh to avoid compounding error from erroneous pose estimations by the model, as shown in Fig. 12.
-0.50-0.250.000.250.500.751.001020Chamfer Distance Error (m)[154][153][196]
Figure 11: Relative object position error of scene-level reconstruction models PhysGen3D [196], Gen3DSR [153], and MIDI [154]. Indicators for min, max, median, and 25th and 75th quantile are shown. We evaluate the difference between the Chamfer distances (CD) between all object pairs in the ground truth scene, and the CDs between the corresponding object pairs in the reconstructed scene with relative object poses estimated by the reconstruction models. Scene reconstruction models tend to overestimate relative object distances, resulting in inaccurate scene reconstructions and potentially impaired scene stability.
SF3D [80]
InstantMesh [27]
One2345 [26]
LGM [96]
Michelangelo [28]
ZeroShape [142]
Real3D [71]
DSO [190]
0.5151050100Reconstruction Time per Sample (s)  0{0}5{5}10{10}15{15}20{20}25{25}Peak VRAM Allocation (GB)